OpenStudy (anonymous):

Let f(x)=x^2+1 and g(x)=3x-5. Find f(g(x)).

OpenStudy (anonymous):

\[f(g(x))=f(3x-5)=(3x-5)^2+1\] then algebra

OpenStudy (anonymous):

g(x) = 3x-5 f(g(x)) = f(3x-5) = (3x-5)^2 +1

OpenStudy (anonymous):

so i just leave my answer as (3x-5)^2+1?

OpenStudy (anonymous):

That should be find or you can leave answer as 9 x^2-30x +26

OpenStudy (anonymous):

if i foil?

OpenStudy (anonymous):

ya if you expand

OpenStudy (anonymous):

what would you recommend i do? this is a written assignment so i cant check whats right or wrong like i can on my hw website.

OpenStudy (anonymous):

then use the quadratic one.

OpenStudy (anonymous):

no points will be deduct if you use quadratic

OpenStudy (anonymous):

if they do deduct, ill blame you haha

OpenStudy (anonymous):

lol, but I am like 99.9 percent sure.

OpenStudy (anonymous):

alright alright